A copper wire of 1mm diameter is stretched by applying a force of 10N. Find the stress in the wire

Grade:11

1 Answers

Vikas TU
7 years ago
Stress  = Force/ Area
Area here it is cross – section part.
Hence,
A = pi*(1/2)^2 = > pi/4 mm^2.
Stress = 10/(pi/4) = > 40/pi = > 12.73 N/mm^2.